The length of a rectangle is 5 centimetres more than twice the width. The perimetre is 82 centimetres
and we have to solve these questions:
1) Find an equasion to represent the perimetre of the rectangle. Write your answer in simplest terms.
2) Solve the equasion to find the width of the rectangle
3) Find the length of the rectangle
This is for my homework which is due TOMORROW!!! i f you could answer there for me it would be GREAT!! thankyou soo much and i have tried to work theese out by going ( l= length) (w= width) l+l+w+w=82cm i dont know if this is right or not! THANKS from izzi

Answer by edjones(8007)

L=2w+5, p=82
.
1) P= 2L+2w = 82
2(2w+5)+2w=82 Substitute 2w+5 for L
4w+10+2w=82
6w+10=82
6w+10-10=82-10
6w=72
6w/6=72/6
w=12 cm Answer 2)
.
3) L=2w+5
L=2*12+5
L=29 cm
.
Ed